A sample of coal powder was analysed by proximate analysis. The observed results were as follows. Moisture content: 43%; volatile content: 28%. If the initial weight of the samples was 0.80 gm, find out the amount of moisture content (in gm), volatile content (in gm), carbon content (in gm) and ash content (in %) content, provided that at the end of the experiment the residual ash was found to be 100.0 mg.


1
Expert's answer
2020-07-09T14:55:21-0400

In Sample of coal powder

Initial wt. of sample was 0.80 gm

Ash is 100 Mg after experiment

So % of ash = 0.1gm*100/0.80

= 12.5 %

Moisture content = 43%

Volatile content = 28%

Amnt. of moisture content = 0.80*43/100 = 0.344gm.

Amnt. of volatile content = 0.80*28/100 = 0.224gm.

Carbon content in gm = 0.80 - (0.344+0.224+0.1)

0.132gm